OpenStudy (anonymous):

Solve the equation by taking square roots: 4(x+2)^2=64

OpenStudy (angela210793):

=2(x+2)=8 x+2=4 x=2

OpenStudy (anonymous):

That's not really right.

OpenStudy (anonymous):

sqrt{4(x+2)^2} = 4(x+2) sqrt{64} = 8, -8 4x + 8 = -8 x = 4 4x+8 = 8 x = 0 x = 0, 4

OpenStudy (angela210793):

cause sqr(x^2)=|x|?

OpenStudy (anonymous):

So I distribute the ^2 ?

OpenStudy (anonymous):

Actually sorry angela, that is correct.

OpenStudy (anonymous):

Wait, no. I was right originally.

OpenStudy (anonymous):

The answer on the answer sheet says -6, 2. Don't over think it.

OpenStudy (angela210793):

C'mon Polpak..make up ur mind :P

OpenStudy (anonymous):

\[4(x+2)^2 = 64\]\[\implies (x+2)^2 =16 \]\[\implies \sqrt{(x+2)^2} = \sqrt{16}\]\[\implies |x+2| = 4\]\[\implies x+2 = 4 \text{ OR } x+2 = -4\] Then solve both of those equations to find what your possible x's are

OpenStudy (anonymous):

Okay, but how do I get -6 ?

OpenStudy (anonymous):

x+2 = -4 Subtract 2 from both sides

OpenStudy (anonymous):

Fist divide both sides by 4, then take the square root of both sides\[(x + 2)^2 = 16\] x+ 2 = 4 and x + 2 = -4 Then finish solving.

OpenStudy (anonymous):

Okay. Thank you so much. I didn't divide by the 4 first, so that's where I got screwed up. Thanks.

OpenStudy (anonymous):

You can always divide by the 2 after you take the square root if you didn't divide by the 4 originally.

OpenStudy (anonymous):

Okay, thank you.